OpenStudy (anonymous):

How to add 4 and 3/5 add with 3 and 4/5?

OpenStudy (anonymous):

What is 4 + 3?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

How does that help?

OpenStudy (anonymous):

What is 3 5ths + 4 5ths?

OpenStudy (anonymous):

can you change 3 and 4/5 to an improper fraction based off what i told you for 4 and 3/5?

OpenStudy (anonymous):

No I dont think so special_g, And polpak 7/5

OpenStudy (anonymous):

Can you rewrite 7/5 as a compound number?

OpenStudy (anonymous):

What is a compound number?

OpenStudy (anonymous):

A number with a whole part and a fractional part. Like 4 3/5

OpenStudy (anonymous):

No I can not,

OpenStudy (anonymous):

7/5 = 5/5 + 2/5 = 1 + 2/5

OpenStudy (anonymous):

Yes I got it, Anything else to this problem? Thank you

OpenStudy (anonymous):

So, what is the answer when you add 4 3/5 plus 3 4/5?

OpenStudy (anonymous):

Well it's 7 and 7/5 which is 7 and 1 and 2/5. So....

OpenStudy (anonymous):

Ok here is my main question, The answer to 4 3/5 plus 3 4/5 is= 7 7/5. So you know how that is a inproper fraction, right? Well how do you make the denaminator bigger than the numinator?

OpenStudy (anonymous):

You divide. \(\frac{7}{5} = 7 \div 5 = 1 \) with remainder 2 = \(1+\frac{2}{5}\)

OpenStudy (anonymous):

Thanks, But I got graded hehe i got 100%, Also do you know any analogies polpak?